# MADE-for-ASD: A Multi-Atlas Deep Ensemble Network for Diagnosing Autism Spectrum Disorder

### v1

@@ Abstract
In response to the global need for efficient early diagnosis of Autism Spectrum Disorder (ASD), this paper bridges the gap between traditional, time-consuming diagnostic methods and potential automated solutions. We propose a multi-atlas deep ensemble network, MADE-for-ASD, that integrates multiple atlases of the brain’s functional magnetic resonance imaging (fMRI) data through a weighted deep ensemble network. Our approach integrates demographic information into the prediction workflow, which enhances ASD diagnosis performance and offers a more holistic perspective on patient profiling. We experiment with the well-known publicly available ABIDE (Autism Brain Imaging Data Exchange) I dataset, consisting of resting state fMRI data from 17 different laboratories around the globe. Our proposed system achieves 75.20% accuracy on the entire dataset and 96.40% on a specific subset – both surpassing reported ASD diagnosis accuracy in ABIDE I fMRI studies. Specifically, our model improves by 4.4 percentage points over prior works on the same amount of data. The model exhibits a sensitivity of 82.90% and a specificity of 69.70% on the entire dataset, and 91.00% and 99.50%, respectively, on the specific subset. We leverage the F-score to pinpoint the top 10 ROI in ASD diagnosis, such as precuneus and anterior cingulate/ventromedial. The proposed system can potentially pave the way for more cost-effective, efficient and scalable strategies in ASD diagnosis. Codes and evaluations are publicly available at TBA.
@@
[a]organization=Curtin University, city=Bentley, postcode=6102, state=WA, country=Australia
@@
[b]organization=BRAC University, city=Dhaka, postcode=1212, country=Bangladesh
@@
[c]organization=Australian National University, city=Canberra, postcode=2601, state=ACT, country=Australia
@@ 4.1 Dataset
We use rs-fMRI and demographic data from ABIDE I [11], the first phase of ABIDE, with 505 autistic individuals and 530 Typical Controls (TCs). We include four key demographic features of the subjects: age (years), sex (male/female), handedness (the dominant hand; left/ambiguous/right) and full-scale IQ (overall intellectual ability) (see Table 6 in B for their distribution in both ASD and TC classes). We also employ a subset of the T1w MRI images from the ABIDE I site, NYU Langone Medical Center, to test our proposed methodology. This subset encompasses 182 subjects, including 78 ASD and 104 TC subjects. We selected this specific subset because it contains the largest amount of data among all participatory sites in ABIDE I. Additionally, previous studies [27, 29] have reported ASD classification performance using the NYU subset, which allows us to compare our model’s performance with theirs. Refer to Table 7 in B for details of the NYU subset having 78 autistic individuals and 104 TC subjects.
@@ 4.2.3 Ablation Study
Although the inclusion of demographic data leads to better performance, these data should be carefully utilised with ML-based practical ASD diagnosis systems so that the system is not biased to particular demographics. Future work should focus on quantifying and mitigating such biases in ML models predicting ASD.

### v2

@@ Abstract
In response to the global need for efficient early diagnosis of Autism Spectrum Disorder (ASD), this paper bridges the gap between traditional, time-consuming diagnostic methods and potential automated solutions. We propose a multi-atlas deep ensemble network, MADE-for-ASD, that integrates multiple atlases of the brain’s functional magnetic resonance imaging (fMRI) data through a weighted deep ensemble network. Our approach integrates demographic information into the prediction workflow, which enhances ASD diagnosis performance and offers a more holistic perspective on patient profiling. We experiment with the well-known publicly available ABIDE (Autism Brain Imaging Data Exchange) I dataset, consisting of resting state fMRI data from 17 different laboratories around the globe. Our proposed system achieves 75.20% accuracy on the entire dataset and 96.40% on a specific subset – both surpassing reported ASD diagnosis accuracy in ABIDE I fMRI studies. Specifically, our model improves by 4.4 percentage points over prior works on the same amount of data. The model exhibits a sensitivity of 82.90% and a specificity of 69.70% on the entire dataset, and 91.00% and 99.50%, respectively, on the specific subset. We leverage the F-score to pinpoint the top 10 ROI in ASD diagnosis, such as precuneus and anterior cingulate/ventromedial. The proposed system can potentially pave the way for more cost-effective, efficient and scalable strategies in ASD diagnosis. Codes and evaluations are publicly available at https://github.com/hasan-rakibul/MADE-for-ASD.
@@
[a]organization=Australian National University, city=Canberra, postcode=2601, state=ACT, country=Australia
@@
[…]
[d]organization=BRAC University, city=Dhaka, postcode=1212, country=Bangladesh
@@ 4.1 Dataset
We use rs-fMRI and demographic data from ABIDE I [11], the first phase of ABIDE, with 505 autistic individuals and 530 Typical Controls (TCs). We include four key demographic features of the subjects: age (years), sex (male/female), handedness (the dominant hand; left/ambiguous/right) and full-scale IQ (overall intellectual ability) (see Table 6 and Table 7 in B for the distribution of demographic information across ASD and TC classes and the number of missing demographic data, respectively.). We also employ a subset of the T1w MRI images from the ABIDE I site, NYU Langone Medical Center, to test our proposed methodology. This subset encompasses 182 subjects, including 78 ASD and 104 TC subjects. We selected this specific subset because it contains the largest amount of data among all participatory sites in ABIDE I. Additionally, previous studies [27, 29] have reported ASD classification performance using the NYU subset, which allows us to compare our model’s performance with theirs. Refer to Table 8 in B for details of the NYU subset having 78 autistic individuals and 104 TC subjects.
@@ 4.2.3 Ablation Study
Although the inclusion of demographic data leads to better performance, these data should be carefully utilised with ML-based practical ASD diagnosis systems so that the system is not biased to particular demographics. Future work should focus on quantifying and mitigating such biases in ML models predicting ASD. Furthermore, exploring demographic differences within the pipeline – such as sex differences in the top ROIs revealed by our algorithm – represents an important avenue for future research. Similarly, evaluating performance on stratified demographic subsets would help ensure that our model performs consistently across different groups.
@@ Appendix B ABIDE I Dataset Statistics
Table 7: Number of missing demographic data in the ABIDE I dataset.
Demographic information # of missing values
Age 0
Sex 0
Handedness 326
Full-scale IQ 72